ELLIPTIC BOUNDARY VALUE PROBLEMS FOR THE INHOMOGENEOUS LAPLACE EQUATION ON BOUNDED

DOMAINS

STEVEN G. KRANTZ SONG-YING LI

Abstract. Elliptic estimates in Hardy classes are proved on domains with minimally smooth boundary. The methodology is different from the original methods of Chang/Krantz/Stein.